40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13291     Entry time: Tue Sep 5 02:07:49 2017     In reply to: 13289     Reply to this: 13300
Author: gautam 
Type: Update 
Category: LSC 
Subject: Low Noise DRMI attempt 

Summary:

Tonight, I was able to lock the DRMI, turn on the whitening filters for the sensing PDs, and also turn on the coil de-whitening filters for ITMX, ITMY and BS. However, I didn't see the expected improvement in the MICH spectrum between ~50-300 Hz sad. Sad.

Details:

I basically went through the list of tasks I made in the previous elog. Some notes:

  • The UGF servos suggested that I had to lower the SRCL gain. I lowered it from -0.055 to -0.025. OLTF measurement using In1/In2 method suggested UGF ~120Hz. I don't know why this should be. Plot to be uploaded later.
  • Since we aren't actuating on the ITMs, I was able to leave their coils de-whitened all the time.
  • For the BS, it was trickier - I had to play around a little with the "Tolerance" setting in Foton while looking at transients (using DTT, not a scope for now) while switching the filters.
  • This transition isn't so robust yet - but eventually I found a setting that worked, and I was able to successfully turn on the de-whitening thrice tonight (but also failed about the same number of times). [GV Oct 6 2017: Remember that the PD whitening has to be turned on for this transition to be successful - otherwise the RMS from the high frequencies saturate the DAC.]
  • The locks were pretty stable. One was ~10mins, one was ~15mins, and I broke both deliberately because I was out of ideas as to why the part of the MICH error signal spectrum that I thought was due to DAC noise didn't improve.
  • I've made a bunch of shell scripts to help with the various switchings - but now that I think of it, I should make these python scripts.

Attachment #1: Comparison of MICH_ERR with and without the BS de-whitening. Note that the two ITMs have their coils de-whitened in both sets of traces.

Attachment #2: Spectra of MICH output and one of the BS coil outputs in both states. The DAC RMS increases by ~30x when the de-whitening is engaged, but is still well within limits.

So it looks like the switching of paths is happening correctly. The "CDS BIO STATUS" MEDM screen also shows the appropriate bits toggling when I turn the de-whitening on/off. There is no broadband coherence with MCF between 50-300 Hz so it seems unlikely that this could be frequency noise.

Clearly I am missing something. But anyways I have a good amount of data, may be useful to put together the post CDS/electronics modification DRMI noise budget. More analysis to follow.

 

Attachment 1: MICH_err_comp.pdf  41 kB  | Hide | Hide all
MICH_err_comp.pdf
Attachment 2: deWhitenedCoil.pdf  65 kB  | Hide | Hide all
deWhitenedCoil.pdf
ELOG V3.1.3-